If cosΦ = C, express in terms of C tan² Φ + 2 cos Φ
1
Expert's answer
2019-03-13T11:16:32-0400

Solution:

"\tan^2\Phi+2\cos\Phi=\frac{\sin^2\Phi} {\cos^2\Phi} +2\cos\Phi=\frac{1-\cos^2\Phi} {\cos^2\Phi} +2\cos\Phi"

Let's take into account that "\cos\Phi=C:"


"\frac{1-\cos^2\Phi} {\cos^2\Phi} +2\cos\Phi=\frac{1-C^2} {C^2} +2C=\frac{1} {C^2}+2C-1"

Answer:

"\frac{1} {C^2}+2C-1"